Lee can frame a cabin in 4 days less than ron When they work together they will do the job in 4 4/5 days how l ong would each of them take to frame the cabin alone.
Answer by checkley77(12844)

X(X-4)/(X+X-4)=4.8
(X^2-4X)/(2X-4)=4.8
X^2-4X=4.8(2X-4)
X^2-4X=9.6X-19.2
X^2-4X-9.6X+19.2=0
X^2-13.6+19.2=0
(X-12)(X-1.6)=0
X-12=0
X=12 ANSWER.
12-4=8 ANSWER.
PROOF:
12*8/12+8=4.8
96/20=4.8
4.8=4.8